A packet of breakfast cereal has the following information on it: Protein = 10.3g per 100g of cereal.

The recommended daily amount of protein for an adult male is 50g. What percentage of the recommended daily amount of protein does 40g of cereal provide?
1
Expert's answer
2013-03-15T08:45:14-0400
One gram of cereal provides

m = 10.3[g]/100 = 0.103[g]

of protein. Therefore, 40g of cereal provide

M = 40m = 4.12[g]

of protein. Therefore, 40g of cereal provide

P = (M/50[g])·100% = (4.12[g]/50[g])·100% = 8.24%

of the recommended daily amount of protein.